Paul’s lawn- mowing service is a profit maximizing, competitive firm. Paul mows lawns for $27 each. His total cost

each day is $280, of which $30 is a fixed cost. He mows 10 lawns a day. What can you say aboutPaul’s short-run decision

regarding shutdown and his long-run decision regarding exit?


1
Expert's answer
2022-03-30T14:14:33-0400

Since price is given as $27, his total revenue, TR, given the quantity of lawns he mows per day (10) equals

$27×10=$270

His total cost, TC, equals $280 and TC=FC+VC

FC=$30, VC=$250

"\pi=\$270-\$280=-\$10"

He makes a loss.


His short run decision would be to decide whether to shutdown or continue in production.

"AC=\frac{TC}{Q}=\frac{\$280}{10}=\$28"

"AVC=\frac{VC}{Q}=\frac{\$250}{10}=\$25"

Since "AC>P>AVC" , i.e.

"\$28>\$27>\$25" , he should keep producing even when he is making a loss.


In the long run, if things do not get better, i.e. he keeps making losses, then he exits the industry.